OpenStudy (anonymous):

Solve the following system of equations for x. - 3x + 7y = 2 - 2x - 3y = - 4 A. x=2/23 B. x=22/23 C. x=-2/23 D. x=-22/23

OpenStudy (anonymous):

What method would you like to use?

OpenStudy (anonymous):

I dont know

OpenStudy (anonymous):

I would use substitution since there is no easy elimination strategy. First solve one equation for one variable.

OpenStudy (anonymous):

-3x+7y=2 y=(2-3x)/7 -2x-3((2-3x)/7)=-4

OpenStudy (anonymous):

you can take it from there